# Characterization of Gaussian Tensor Ensembles

### v1

@@ 2.1 Warm-up : proof of the original theorem for vectors
First, the law of a Gaussian vector belonging to $\mathcal{N}(0,\sigma^{2}I_{N})=\mathcal{N}(0,\sigma^{2})^{\otimes N}$, having for density $H\mapsto\exp(\frac{1}{2}|H|_{2}^{2}-\frac{N}{2}\ln(2\pi\sigma^{2}))$, is product and rotationally invariant. Reciprocally, let $N\geq 2$ and $\mu$ be a law on $\mathbb{R}^{N}$ product and rotationally invariant. Assume in a first time that $\mu$ has a smooth density $f:\mathbb{R}^{N}\rightarrow(0,\infty)$. As $\mu$ is product, $f$ must be of the form $f(H)=\prod_{i=1}^{N}f_{i}(H_{i})$, that is
@@
Hence, taking the derivative along one coordinate we get
$\partial_{i}\ln f(H)=g^{\prime}_{i}(H_{i}),$
so $\partial_{i}\ln f(H)$ must depend only on $H_{i}$. Moreover, the rotational invariance implies $\ln f(H)=\psi(|H|^{2}_{2})$, so we have
$\partial_{i}\ln f(H)=2\psi^{\prime}(|H|^{2}_{2})H_{i}.$
@@ 2.2 Proof of the Main Theorem
$A=\frac{\partial U_{\theta}^{T}}{\partial\theta}U_{\theta}=\begin{pmatrix}0&-1&\cdots&0\\ 1&0&\cdots&0\\ \cdots&\cdots&\ddots&\cdots\\ 0&\cdots&\cdots&0\\ \end{pmatrix}.$
@@
$\left(\frac{\partial H}{\partial\theta}\right)_{i_{1},\ldots,i_{p}}=\sum_{k=1}^{N}(A_{i_{1}k}H_{k,i_{2},\ldots,i_{p}}+A_{i_{2}k}H_{i_{1},k,\ldots,i_{p}}+\ldots+A_{i_{p}k}H_{i_{1},\ldots,i_{p-1},k}).$
Lemma 1.
Let $g_{1},g_{2},g_{3}$ be three continuous and differentiable functions such that
$g_{1}(xy)=g_{2}(x)+g_{3}(y).$
Then, there exists $\alpha,\beta_{1},\beta_{2},\beta_{3}$ such that for all $k\in\{1,2,3\}$,
$g_{k}(u)=\alpha\ln u+\beta_{k},$
with $\beta_{1}=\beta_{2}+\beta_{3}$.
Proof of the Theorem..
@@
$f(H)=\prod_{i_{1}\leq\ldots\leq i_{p}}f_{i_{1},\ldots,i_{p}}(H_{i_{1},\ldots,i_{p}}).$
By orthogonal invariance,
$\frac{\partial}{\partial\theta}\ln f(H)=0.$
Then using the computation of Proposition 1, we find
00 0 $=\sum_{i_{1}\leq\ldots\leq i_{p}}\frac{1}{f_{i_{1},\ldots,i_{p}}(H_{i_{1},\ldots,i_{p}})}\frac{\partial f_{i_{1},\ldots,i_{p}}}{\partial H_{i_{1},\ldots,i_{p}}}\frac{\partial H_{i_{1},\ldots,i_{p}}}{\partial\theta}$
$=\sum_{i_{1}\leq\ldots\leq i_{p}}\frac{1}{f_{i_{1},\ldots,i_{p}}}\frac{\partial f_{i_{1},\ldots,i_{p}}}{\partial H_{i_{1},\ldots,i_{p}}}\sum_{k=1}^{N}(A_{i_{1}k}H_{k,i_{2},\ldots,i_{p}}+A_{i_{2}k}H_{i_{1},k,\ldots,i_{p}}+\ldots+A_{i_{p}k}H_{i_{1},\ldots,i_{p-1},k}).$
Moreover, we know that $A_{i,k}$ is equal to $-1$ if $(i,k)=(1,2)$, to $1$ if $(i,k)=(2,1)$ and to 00 otherwise. We introduce a brief notation for ease of notation in the sums with multi-indices. For given integers $r\in\{1,\ldots,p-1\}$ and $i_{r+1},\ldots,i_{p}$, let us denote $f_{a_{1},\ldots,a_{r}}$ for $f_{a_{1},\ldots,a_{r},i_{p-r},\ldots,i_{p}}$ and $H_{a_{1},\ldots,a_{r}}$ for $H_{a_{1},\ldots,a_{r},i_{p-r},\ldots,i_{p}}$. In particular, $f_{1}(H_{1})=f_{1,i_{2},\ldots,i_{p}}(H_{1,i_{2},\ldots,i_{p}})$. The previous equations gives finally
00 0 $=\sum_{3\leq i_{2}\leq\ldots\leq i_{p}}\left[\frac{-1}{f_{1}(H_{1})}\frac{\partial f_{1}}{\partial H_{1}}H_{2}+\frac{1}{f_{2}(H_{2})}\frac{\partial f_{2}}{\partial H_{2}}H_{1}\right]$ (A)
$+\sum_{3\leq i_{3}\leq\ldots\leq i_{p}}\left[2\left(\frac{-1}{f_{1,1}(H_{1,1})}\frac{\partial f_{1,1}}{\partial H_{1,1}}+\frac{1}{f_{2,2}(H_{2,2})}\frac{\partial f_{2,2}}{\partial H_{2,2}}\right)H_{1,2}+\frac{1}{f_{1,2}(H_{1,2})}\frac{\partial f_{1,2}}{\partial H_{1,2}}(H_{1,1}-H_{2,2})\right]$
$+\sum_{3\leq i_{4}\leq\ldots\leq i_{p}}\left[\ldots\right]$
The different brackets depend on mutually exclusive sets of variables, and their sum is 00, so each must be a constant and in particular for given integers $i_{p}\geq\ldots\geq i_{2}\geq 3$, there exists $C=C_{i_{2},\ldots,i_{p}}$ such that
@@
and now applying Lemma 1 we find that the constant $C$ must be zero. That means
@@
and this must be a constant $c^{\prime}=c^{\prime}_{i_{2},\ldots,i_{p}}$, which gives
@@
After integrating we find $\ln f_{1}(H_{1})=\frac{c^{\prime}}{2}H_{1}^{2}$, so $c^{\prime}$ must be negative ($f_{1}$ is a density) and we denote it $-2a$ in place of, with $a\in\mathbb{R}^{*}_{+}$ (depending on ${i_{2},\ldots,i_{p}}$). Finally, we conclude from all of this
$f_{1,i_{2},\ldots,i_{p}}(H_{1,i_{2},\ldots,i_{p}})=\exp(-aH_{1,i_{2},\ldots,i_{p}}^{2}).$
The same holds for $f_{2}(H_{2})$ giving $f_{2,i_{2},\ldots,i_{p}}(H_{2,i_{2},\ldots,i_{p}})=\exp(-aH_{2,i_{2},\ldots,i_{p}}^{2})$. Note that involving the symmetries and orthogonal matrices $U_{\theta}$ with the rotation acting at different positions than $(1,2)$, it follows that any $p$-tuple of indices with at least one index $i_{k}$ distinct from all other ones gives a contribution $f_{i_{1},\ldots,i_{p}}(H_{i_{1},\ldots,i_{p}})=\exp(-a_{i_{1},\ldots,i_{k-1},i_{k+1},\ldots,i_{p}}H_{i_{1},\ldots,i_{p}}^{2})$. Now, the second bracket in equation $(A)$ gives for $i_{p}\geq\ldots\geq i_{3}\geq 3$,
$2\left(\frac{-1}{f_{1,1}(H_{1,1})}\frac{\partial f_{1,1}}{\partial H_{1,1}}+\frac{1}{f_{2,2}(H_{2,2})}\frac{\partial f_{2,2}}{\partial H_{2,2}}\right)-2a_{1,2,i_{3},\ldots,i_{p}}(H_{1,1}-H_{2,2})=\frac{C_{i_{3},\ldots,i_{p}}}{H_{1,2}}=0,$
@@
$\frac{1}{f_{1,1}(H_{1,1})}\frac{\partial f_{1,1}}{\partial H_{1,1}}+aH_{1,1}=\frac{1}{f_{2,2}(H_{2,2})}\frac{\partial f_{2,2}}{\partial H_{2,2}}+aH_{2,2}=b_{i_{3},\ldots,i_{p}},$
that is finally
$f_{1,1,i_{3},\ldots,i_{p}}(H_{1,1,i_{3},\ldots,i_{p}})=\exp(-\frac{a}{2}H_{1,1,i_{3},\ldots,i_{p}}^{2}-bH_{1,1,i_{3},\ldots,i_{p}}).$
But as all invariants are expressible in terms of the trace invariants, the two only ones which can appear are the melon (corresponding to the Frobenius norm) and the bouquet (corresponding to the paired trace), and the constants must satisfy $a_{i_{1},\ldots,i_{p}}=\frac{a_{1,\ldots,p}}{\Gamma_{i_{1},\ldots,i_{p}}}$, $b_{i_{1},i_{1},\ldots,i_{p/2},i_{p/2}}=\frac{b}{\Gamma_{i_{1},i_{1},\ldots,i_{p/2},i_{p/2}}}$ and $b_{i_{1},\ldots,i_{p}}=0$ otherwise, leading to
$f(H)$ $=\exp\left(-\sum_{i_{1}\leq\ldots\leq i_{p}}\frac{a}{\Gamma_{i_{1},\ldots,i_{p}}}H_{i_{1},\ldots,i_{p}}^{2}+\mathds{1}_{p\text{ even}}\sum_{i_{1}\leq\ldots\leq i_{p/2}}\frac{b}{\Gamma_{i_{1},i_{1},\ldots,i_{p/2},i_{p/2}}}H_{i_{1},i_{1},\ldots,i_{p/2},i_{p/2}}+c\right)$
$=\exp\left(-a\sum_{i_{1},\ldots,i_{p}}H_{i_{1},\ldots,i_{p}}^{2}+b\mathds{1}_{p\text{ even}}\sum_{i_{1},\ldots,i_{p/2}}H_{i_{1},i_{1},\ldots,i_{p/2},i_{p/2}}+c\right).$
@@
$\sum_{k=1}^{N}U_{i_{1}k}H_{k,i_{2},\ldots,i_{p}}$ $=\sum_{j_{1},\ldots,j_{p}}H^{\prime}_{j_{1},\ldots,j_{p}}\prod_{t=2}^{p}U_{j_{t}i_{t}}\sum_{k=1}^{N}U_{i_{1}k}U^{T}_{k,j_{1}}$
$=\sum_{j_{1},\ldots,j_{p}}H^{\prime}_{j_{1},\ldots,j_{p}}\prod_{t=2}^{p}U_{j_{t}i_{t}}\times\mathbf{1}_{j_{1}=i_{1}},$
@@
$\sum_{k=1}^{N}U_{i_{1}k}H_{k,i_{2},\ldots,i_{p}}=\sum_{j_{2},\ldots,j_{p}}H^{\prime}_{i_{1},j_{2},\ldots,j_{p}}\prod_{t=2}^{p}U_{j_{t}i_{t}}$ (1)
@@
$\left(\frac{\partial H}{\partial\theta}\right)_{i_{1},\ldots,i_{p}}$ $=\sum_{j_{1},\ldots,j_{p}}H^{\prime}_{j_{1},\ldots,j_{p}}\sum_{r=1}^{p}\prod_{t=1\atop t\neq r}^{p}U_{j_{t}i_{t}}\left(\frac{\partial U_{j_{r},i_{r}}}{\partial\theta}\right)$
$=\sum_{r=1}^{p}\sum_{j_{r}}\left(\frac{\partial U_{j_{r},i_{r}}}{\partial\theta}\right)\sum_{j_{1},\ldots,j_{p}\atop\widehat{j_{r}}}H^{\prime}_{j_{1},\ldots,j_{p}}\prod_{t=1\atop t\neq r}^{p}U_{j_{t}i_{t}}$
$=\sum_{r=1}^{p}\sum_{k=1}^{N}\left(\sum_{j_{r}}\left(\frac{\partial U}{\partial\theta}\right)^{T}_{i_{r},j_{r}}\times U_{j_{r}k}\right)H_{i_{1},\ldots,i_{r-1},k,i_{r+1},\ldots,i_{p}}$
where we used equation 1 for the last line. Replacing by $A$, we finally get the result
$\left(\frac{\partial H}{\partial\theta}\right)_{i_{1},\ldots,i_{p}}=\sum_{k=1}^{N}(A_{i_{1}k}H_{k,i_{2},\ldots,i_{p}}+A_{i_{2}k}H_{i_{1},k,\ldots,i_{p}}+\ldots+A_{i_{p}k}H_{i_{1},\ldots,i_{p-1},k}).$
Proof of Lemma 1.
Differentiating the relation $g_{1}(xy)=g_{2}(x)+g_{3}(y)$ with respect to $x$ gives
$g^{\prime}_{1}(xy)=yg^{\prime}_{2}(x),$
and after integrating with respect to $y$, we get
$g_{1}(xy)=xg^{\prime}_{2}(x)\ln y+h(x)$
Hence, inserting this equation in the initial relation, we have
@@
The right-hand side term must be independent of $y$ which means that there exists constants $a,b_{3}$ such that $xg^{\prime}_{2}(x)=a$ and $h(x)-g_{2}(x)=b_{3}$, that is
$g_{3}(y)=a\ln y+b_{3},$
and by integration of $g^{\prime}_{2}(x)=a/x$ we also have
$g_{2}(x)=a\ln x+b_{2}.$
Finally, this gives $g_{1}(xy)=a\ln(xy)+b_{2}+b_{3}$. ∎
Unitary and symplectic cases.
We adapt the proof for hermitian and self-dual hermitian tensors. For $\theta\in[0,2\pi]$, we take the same matrix $U_{\theta}$ which is both orthogonal, unitary and symplectic (of size $2N$ in this last case), that is
$U_{\theta}=\begin{pmatrix}\cos{\theta}&\sin{\theta}&0&\cdots&0\\ -\sin{\theta}&\cos\theta&0&\cdots&0\\ 0&\cdots&1&\cdots&0\\ 0&\cdots&\cdots&\ddots&0\\ 0&\cdots&0&\cdots&1\\ \end{pmatrix}\quad\mathrm{and}\quad A=\frac{\partial U_{\theta}^{T}}{\partial\theta}U_{\theta}=\begin{pmatrix}0&-1&\cdots&0\\ 1&0&\cdots&0\\ 0&\cdots&\ddots&0\\ 0&\cdots&\cdots&0\\ \end{pmatrix},$
@@
$U_{\theta}=\begin{pmatrix}\mathbf{1}\cos{\theta}-\mathbf{e_{2}}\sin{\theta}&0&\cdots&0\\ 0&\mathbf{1}&\cdots&0\\ \cdots&\cdots&\ddots&\cdots\\ 0&\cdots&\cdots&\mathbf{1}\\ \end{pmatrix}\quad\mathrm{and}\quad A=\begin{pmatrix}\mathbf{e_{2}}&0&\cdots&0\\ 0&0&\cdots&0\\ \cdots&\cdots&\cdots&\cdots\\ 0&\cdots&\cdots&0\\ \end{pmatrix}.$
The computation of the derivatives of $H$ with respect to $\theta$ is exactly the same. Now recall that if $H$ is hermitian (resp. $\mathrm{H}$ is self-dual hermitian identified as $Q$ with $(p/2)$-quaternionic entries), it is of the form $H=H^{(0)}+iH^{(1)}$ (resp. $Q=\sum_{\epsilon\in\mathcal{E}}Q^{(\epsilon)}\otimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}}$) with $H^{(0)}\in\mathcal{S}^{(p)}(N),H^{(1)}\in\mathcal{A}^{(p)}(N)$ (resp. $Q^{(\epsilon)}\in\mathcal{S}^{(p)}(N)$ if $\epsilon\in\mathcal{E}$, $Q^{(\epsilon)}\in\mathcal{A}^{(p)}(N)$ otherwise). Hence if $H$, resp. $\mathrm{H}$, is a tensor hermitian, resp. self-dual hermitian, with independent entries and a law unitary, resp. symplectic, invariant, their laws have densities (without loss of generality after adding a tensor from the GUTE, resp. GSTE) of the form
$f(H)=\prod_{i_{1}\leq\ldots\leq i_{p}}f_{i_{1},\ldots,i_{p}}^{(0)}(H^{(0)}_{i_{1},\ldots,i_{p}})\prod_{i_{1}\leq\ldots\leq i_{p}\atop\text{not paired}}f_{i_{1},\ldots,i_{p}}^{(1)}(H^{(1)}_{i_{1},\ldots,i_{p}}),$
@@
$f(\mathrm{H})=\prod_{\epsilon\in\mathcal{E}}\prod_{i_{1}\leq\ldots\leq i_{p}}f_{i_{1},\ldots,i_{p}}^{(\epsilon)}(Q^{(\epsilon)}_{i_{1},\ldots,i_{p}})\prod_{\epsilon\notin\mathcal{E}}\prod_{i_{1}\leq\ldots\leq i_{p}\atop\text{not paired}}f_{i_{1},\ldots,i_{p}}^{(\epsilon)}(Q^{(\epsilon)}_{i_{1},\ldots,i_{p}}).$
Then, the invariance under respective similarity transformations gives
00 0 $=\frac{\partial}{\partial\theta}\ln f(H)$
$=\left[\sum_{i_{1}\leq\ldots\leq i_{p}}\frac{1}{f^{(0)}_{i_{1},\ldots,i_{p}}(H^{(0)}_{i_{1},\ldots,i_{p}})}\frac{\partial f^{(0)}_{i_{1},\ldots,i_{p}}}{\partial H^{(0)}_{i_{1},\ldots,i_{p}}}\frac{\partial H^{(0)}_{i_{1},\ldots,i_{p}}}{\partial\theta}\right]+\left[\sum_{i_{1}\leq\ldots\leq i_{p}\atop\text{not paired}}\frac{1}{f^{(1)}_{i_{1},\ldots,i_{p}}(H^{(1)}_{i_{1},\ldots,i_{p}})}\frac{\partial f^{(1)}_{i_{1},\ldots,i_{p}}}{\partial H^{(1)}_{i_{1},\ldots,i_{p}}}\frac{\partial H^{(1)}_{i_{1},\ldots,i_{p}}}{\partial\theta}\right],$
@@
$0=\sum_{\epsilon\in\mathcal{E}}\left[\sum_{i_{1}\leq\ldots\leq i_{p}}\frac{1}{f^{(\epsilon)}_{i_{1},\ldots,i_{p}}(Q^{(\epsilon)}_{i_{1},\ldots,i_{p}})}\frac{\partial f^{(\epsilon)}_{i_{1},\ldots,i_{p}}}{\partial Q^{(\epsilon)}_{i_{1},\ldots,i_{p}}}\frac{\partial Q^{(\epsilon)}_{i_{1},\ldots,i_{p}}}{\partial\theta}\right]+\sum_{\epsilon\notin\mathcal{E}}\left[\sum_{i_{1}\leq\ldots\leq i_{p}\atop\text{not paired}}\frac{1}{f^{(\epsilon)}_{i_{1},\ldots,i_{p}}(Q^{(\epsilon)}_{i_{1},\ldots,i_{p}})}\frac{\partial f^{(\epsilon)}_{i_{1},\ldots,i_{p}}}{\partial Q^{(\epsilon)}_{i_{1},\ldots,i_{p}}}\frac{\partial Q^{(\epsilon)}_{i_{1},\ldots,i_{p}}}{\partial\theta}\right].$

### v2

@@ 2.1 Warm-up: Proof of the original theorem for vectors
First, the law of a Gaussian vector belonging to $\mathcal{N}\bigl(0,\sigma^{2}I_{N}\bigr)=\mathcal{N}\bigl(0,\sigma^{2}\bigr)^{\otimes N}$, having for density $H\mapsto\exp\bigl(\frac{1}{2}|H|_{2}^{2}-\frac{N}{2}\ln\bigl(2\pi\sigma^{2}\bigr)\bigr)$, is product and rotationally invariant. Reciprocally, let $N\geq 2$ and $\mu$ be a law on $\mathbb{R}^{N}$ product and rotationally invariant. Assume in a first time that $\mu$ has a smooth density $f\colon\mathbb{R}^{N}\mapsto(0,\infty)$. As $\mu$ is product, $f$ must be of the form $f(H)=\prod_{i=1}^{N}f_{i}(H_{i})$, that is,
@@
Hence, taking the derivative along one coordinate we get $\partial_{i}\ln f(H)=g^{\prime}_{i}(H_{i})$, so $\partial_{i}\ln f(H)$ must depend only on $H_{i}$. Moreover, the rotational invariance implies $\ln f(H)=\psi\bigl(|H|^{2}_{2}\bigr)$, so we have
$\partial_{i}\ln f(H)=2\psi^{\prime}\bigl(|H|^{2}_{2}\bigr)H_{i}.$
@@ 2.2 Proof of the main theorem
$A=\frac{\partial U_{\theta}^{\mathsf{T}}}{\partial\theta}U_{\theta}=\begin{pmatrix}0&-1&\cdots&0\\ 1&0&\cdots&0\\ \cdots&\cdots&\ddots&\cdots\\ 0&\cdots&\cdots&0\\ \end{pmatrix}$.
@@
$\left(\frac{\partial H}{\partial\theta}\right)_{i_{1},\dots,i_{p}}=\sum_{k=1}^{N}(A_{i_{1}k}H_{k,i_{2},\dots,i_{p}}+A_{i_{2}k}H_{i_{1},k,\dots,i_{p}}+\dots+A_{i_{p}k}H_{i_{1},\dots,i_{p-1},k}).$
Lemma 2.2.
Let $g_{1}$, $g_{2}$, $g_{3}$ be three continuous and differentiable functions such that $g_{1}(xy)=g_{2}(x)+g_{3}(y)$. Then, there exists $\alpha$, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$ such that for all $k\in\{1,2,3\}$, $g_{k}(u)=\alpha\ln u+\beta_{k}$, with $\beta_{1}=\beta_{2}+\beta_{3}$.
Proof of Theorem 1.8..
@@
$f(H)=\prod_{i_{1}\leq\dots\leq i_{p}}f_{i_{1},\dots,i_{p}}(H_{i_{1},\dots,i_{p}}).$
By orthogonal invariance, $\frac{\partial}{\partial\theta}\ln f(H)=0$. Then using the computation of Proposition 2.1, we find
$0$ $=\sum_{i_{1}\leq\dots\leq i_{p}}\frac{1}{f_{i_{1},\dots,i_{p}}(H_{i_{1},\dots,i_{p}})}\frac{\partial f_{i_{1},\dots,i_{p}}}{\partial H_{i_{1},\dots,i_{p}}}\frac{\partial H_{i_{1},\dots,i_{p}}}{\partial\theta}$
$=\sum_{i_{1}\leq\dots\leq i_{p}}\frac{1}{f_{i_{1},\dots,i_{p}}}\frac{\partial f_{i_{1},\dots,i_{p}}}{\partial H_{i_{1},\dots,i_{p}}}\sum_{k=1}^{N}(A_{i_{1}k}H_{k,i_{2},\dots,i_{p}}+A_{i_{2}k}H_{i_{1},k,\dots,i_{p}}+\dots+A_{i_{p}k}H_{i_{1},\dots,i_{p-1},k}).$
Moreover, we know that $A_{i,k}$ is equal to 1-1 if $(i,k)=(1,2)$, to $1$ if $(i,k)=(2,1)$ and to 0 otherwise. We introduce a brief notation for ease of notation in the sums with multi-indices. For given integers $r\in\{1,\dots,p-1\}$ and $i_{r+1},\dots,i_{p}$, let us denote $f_{a_{1},\dots,a_{r}}$ for $f_{a_{1},\dots,a_{r},i_{p-r},\dots,i_{p}}$ and $H_{a_{1},\dots,a_{r}}$ for $H_{a_{1},\dots,a_{r},i_{p-r},\dots,i_{p}}$. In particular, $f_{1}(H_{1})=f_{1,i_{2},\dots,i_{p}}(H_{1,i_{2},\dots,i_{p}})$. The previous equations gives finally
$0={}$ $\sum_{3\leq i_{2}\leq\dots\leq i_{p}}\left[\frac{-1}{f_{1}(H_{1})}\frac{\partial f_{1}}{\partial H_{1}}H_{2}+\frac{1}{f_{2}(H_{2})}\frac{\partial f_{2}}{\partial H_{2}}H_{1}\right]$
$+\sum_{3\leq i_{3}\leq\dots\leq i_{p}}\left[2\left(\frac{-1}{f_{1,1}(H_{1,1})}\frac{\partial f_{1,1}}{\partial H_{1,1}}+\frac{1}{f_{2,2}(H_{2,2})}\frac{\partial f_{2,2}}{\partial H_{2,2}}\right)H_{1,2}\right$.
$\left.+\frac{1}{f_{1,2}(H_{1,2})}\frac{\partial f_{1,2}}{\partial H_{1,2}}(H_{1,1}-H_{2,2})\right]+\sum_{3\leq i_{4}\leq\dots\leq i_{p}}\left[\dots\right].$ (2.1)
The different brackets depend on mutually exclusive sets of variables, and their sum is 0, so each must be a constant and in particular for given integers $i_{p}\geq\dots\geq i_{2}\geq 3$, there exists $C=C_{i_{2},\dots,i_{p}}$} such that
@@
and now applying Lemma 2.2 we find that the constant $C$ must be zero. That means
@@
and this must be a constant $c^{\prime}=c^{\prime}_{i_{2},\dots,i_{p}}$, which gives
@@
After integrating we find $\ln f_{1}(H_{1})=\frac{c^{\prime}}{2}H_{1}^{2}$, so $c^{\prime}$ must be negative ($f_{1}$ is a density) and we denote it 2a-2a in place of, with $a\in\mathbb{R}^{*}_{+}$ (depending on $i_{2},\dots,i_{p}$}). Finally, we conclude from all of this $f_{1,i_{2},\dots,i_{p}}(H_{1,i_{2},\dots,i_{p}})=\exp\bigl(-aH_{1,i_{2},\dots,i_{p}}^{2}\bigr)$. The same holds for $f_{2}(H_{2})$ giving $f_{2,i_{2},\dots,i_{p}}(H_{2,i_{2},\dots,i_{p}})=\exp\bigl(-aH_{2,i_{2},\dots,i_{p}}^{2}\bigr)$. Note that involving the symmetries and orthogonal matrices $U_{\theta}$ with the rotation acting at different positions than $(1,2)$, it follows that any $p$-tuple of indices with at least one index $i_{k}$ distinct from all other ones gives a contribution $f_{i_{1},\dots,i_{p}}(H_{i_{1},\dots,i_{p}})=\exp\bigl(-a_{i_{1},\dots,i_{k-1},i_{k+1},\dots,i_{p}}H_{i_{1},\dots,i_{p}}^{2}\bigr)$}. Now, the second bracket in equation (2.1) gives for $i_{p}\geq\dots\geq i_{3}\geq 3$,
$2\left(\frac{-1}{f_{1,1}(H_{1,1})}\frac{\partial f_{1,1}}{\partial H_{1,1}}+\frac{1}{f_{2,2}(H_{2,2})}\frac{\partial f_{2,2}}{\partial H_{2,2}}\right)-2a_{1,2,i_{3},\dots,i_{p}}(H_{1,1}-H_{2,2})=\frac{C_{i_{3},\dots,i_{p}}}{H_{1,2}}=0,$
@@
$\frac{1}{f_{1,1}(H_{1,1})}\frac{\partial f_{1,1}}{\partial H_{1,1}}+aH_{1,1}=\frac{1}{f_{2,2}(H_{2,2})}\frac{\partial f_{2,2}}{\partial H_{2,2}}+aH_{2,2}=b_{i_{3},\dots,i_{p}},$
that is, finally
$f_{1,1,i_{3},\dots,i_{p}}(H_{1,1,i_{3},\dots,i_{p}})=\exp\left(-\frac{a}{2}H_{1,1,i_{3},\dots,i_{p}}^{2}-bH_{1,1,i_{3},\dots,i_{p}}\right).$
But as all invariants are expressible in terms of the trace invariants, the two only ones which can appear are the melon (corresponding to the Frobenius norm) and the bouquet (corresponding to the paired trace), and the constants must satisfy $a_{i_{1},\dots,i_{p}}=\frac{a_{1,\dots,p}}{\Gamma_{i_{1},\dots,i_{p}}}$, $b_{i_{1},i_{1},\dots,i_{p/2},i_{p/2}}=\smash{\frac{b}{\Gamma_{i_{1},i_{1},\dots,i_{p/2},i_{p/2}}}}$ and $b_{i_{1},\dots,i_{p}}=0$ otherwise, leading to
$f(H)={}$ $\exp\Bigg(-\sum_{i_{1}\leq\dots\leq i_{p}}\frac{a}{\Gamma_{i_{1},\dots,i_{p}}}H_{i_{1},\dots,i_{p}}^{2}$
$+\mathds{1}_{p\text{ even}}\sum_{i_{1}\leq\dots\leq i_{p/2}}\frac{b}{\Gamma_{i_{1},i_{1},\dots,i_{p/2},i_{p/2}}}H_{i_{1},i_{1},\dots,i_{p/2},i_{p/2}}+c\Bigg)$
$={}$ $\exp\Bigg(-a\sum_{i_{1},\dots,i_{p}}H_{i_{1},\dots,i_{p}}^{2}+b\mathds{1}_{p\text{ even}}\sum_{i_{1},\dots,i_{p/2}}H_{i_{1},i_{1},\dots,i_{p/2},i_{p/2}}+c\Bigg).$
@@
$\sum_{k=1}^{N}U_{i_{1}k}H_{k,i_{2},\dots,i_{p}}$ $=\sum_{j_{1},\dots,j_{p}}H^{\prime}_{j_{1},\dots,j_{p}}\prod_{t=2}^{p}U_{j_{t}i_{t}}\sum_{k=1}^{N}U_{i_{1}k}U^{\mathsf{T}}_{k,j_{1}}$
$=\sum_{j_{1},\dots,j_{p}}H^{\prime}_{j_{1},\dots,j_{p}}\prod_{t=2}^{p}U_{j_{t}i_{t}}\times\mathbf{1}_{j_{1}=i_{1}},$
@@
$\sum_{k=1}^{N}U_{i_{1}k}H_{k,i_{2},\dots,i_{p}}=\sum_{j_{2},\dots,j_{p}}H^{\prime}_{i_{1},j_{2},\dots,j_{p}}\prod_{t=2}^{p}U_{j_{t}i_{t}}$ (2.2)
@@
[…]
$=\sum_{r=1}^{p}\sum_{k=1}^{N}\left(\sum_{j_{r}}\left(\frac{\partial U}{\partial\theta}\right)^{\mathsf{T}}_{i_{r},j_{r}}\times U_{j_{r}k}\right)H_{i_{1},\dots,i_{r-1},k,i_{r+1},\dots,i_{p}},$
where we used equation (2.2) for the last line. Replacing by $A$, we finally get the result
$\left(\frac{\partial H}{\partial\theta}\right)_{i_{1},\dots,i_{p}}=\sum_{k=1}^{N}(A_{i_{1}k}H_{k,i_{2},\dots,i_{p}}+A_{i_{2}k}H_{i_{1},k,\dots,i_{p}}+\dots+A_{i_{p}k}H_{i_{1},\dots,i_{p-1},k}).$
Proof of Lemma 2.2.
Differentiating the relation $g_{1}(xy)=g_{2}(x)+g_{3}(y)$ with respect to $x$ gives $g^{\prime}_{1}(xy)=yg^{\prime}_{2}(x)$, and after integrating with respect to $y$, we get $g_{1}(xy)=xg^{\prime}_{2}(x)\ln y+h(x)$ Hence, inserting this equation in the initial relation, we have
@@
[…]
Unitary and symplectic cases. We adapt the proof for hermitian and self-dual hermitian tensors. For $\theta\in[0,2\pi]$, we take the same matrix $U_{\theta}$ which is both orthogonal, unitary and symplectic (of size $2N$ in this last case), that is,
[…]
@@
[…]
The computation of the derivatives of $H$ with respect to $\theta$ is exactly the same. Now recall that if $H$ is hermitian (resp. $\mathrm{H}$ is self-dual hermitian identified as $Q$ with $(p/2)$-quaternionic entries), it is of the form $H=H^{(0)}+{\rm i}H^{(1)}$ (resp. $Q=\sum_{\epsilon\in\mathcal{E}}Q^{(\epsilon)}\bigotimes_{s=1}^{p/2}\mathbf{e_{\epsilon_{s}}}$) with $H^{(0)}\in\mathcal{S}^{(p)}(N)$, $H^{(1)}\in\mathcal{A}^{(p)}(N)$ (resp. $Q^{(\epsilon)}\in\mathcal{S}^{(p)}(N)$ if $\epsilon\in\mathcal{E}$, $Q^{(\epsilon)}\in\mathcal{A}^{(p)}(N)$ otherwise). Hence if $H$, resp. $\mathrm{H}$, is a tensor hermitian, resp. self-dual hermitian, with independent entries and a law unitary, resp. symplectic, invariant, their laws have densities (without loss of generality after adding a tensor from the GUTE, resp. GSTE) of the form
[…]
@@
[…]
Denote $\mathbf{i}=(i_{1},\dots,i_{p})$. Then, the invariance under respective similarity transformations gives
$0$ $=\frac{\partial}{\partial\theta}\ln f(H)$
[…]
@@
[…]